consider matrix A=[101 212 313 111] find the nullity and rank


1
Expert's answer
2021-09-05T18:57:17-0400
"A=\begin{bmatrix}
 1 & 0 & 1 \\
 2 & 1 & 2 \\
 3 & 1 & 3 \\
 1 & 1 & 1 \\
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 3 & 1 & 3 \\
 1 & 1 & 1 \\
\end{bmatrix}"

"R_3=R_3-3R_1"


"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 1 & 0 \\
 1 & 1 & 1 \\
\end{bmatrix}"

"R_4=R_4-R_1"


"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 1 & 0 \\
 0 & 1 & 0 \\
\end{bmatrix}"

"R_3=R_3-R_2"


"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
 0 & 1 & 0 \\
\end{bmatrix}"

"R_4=R_4-R_2"


"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
 0 & 0 & 0 \\
\end{bmatrix}"

The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is "2."


In the case of an "m\times n" matrix, the dimension of the domain is "n," the number of columns in the matrix. 

By the Rank-Nullity Theorem


"{\displaystyle \operatorname {Rank} (A)+\operatorname {Nullity} (A)=n.}"

"{Nullity} (A)=3-2=1"

The nullity of the given matrix "A" is "1."

The rank of the given matrix "A" is "2."